# A Clinically-Grounded Two-Stage Framework for Renal CT Report Generation

### v1

@@ Abstract
Generating radiology reports from CT scans remains a complex task due to the nuanced nature of medical imaging and the variability in clinical documentation. In this study, we propose a two-stage framework for generating renal radiology reports from 2D CT slices. First, we extract structured abnormality features using a multi-task learning model trained to identify lesion attributes such as location, size, enhancement, and attenuation. These extracted features are subsequently combined with the corresponding CT image and fed into a fine-tuned vision-language model to generate natural language report sentences aligned with clinical findings. We conduct experiments on a curated dataset of renal CT studies with manually annotated sentence–slice–feature triplets and evaluate performance using both classification metrics and natural language generation metrics. Our results demonstrate that the proposed model outperforms random baselines across all abnormality types, and the generated reports capture key clinical content with reasonable textual accuracy. This exploratory work highlights the feasibility of modular, feature-informed report generation for renal imaging. Future efforts will focus on extending this pipeline to 3D CT volumes and further improving clinical fidelity in multimodal medical AI systems.
@@ 1 Introduction
Renal cancer, also known as kidney cancer, is a disease with a rising global incidence, with an estimated 400,000 new cases diagnosed annually and approximately 175,000 deaths worldwide each year [1]. Current projections indicate that the incidence will continue to rise over the next decade, underscoring the urgency of addressing this growing public health challenge. In the United States alone, more than 81,000 new cases and over 14,000 deaths are expected in 2024 [2]. Although mortality rates have declined by 1–2% annually in recent years due to advances in treatment, the overall incidence continues to increase [2]. The widespread adoption of cross-sectional imaging, particularly computed tomography (CT), has contributed to an increase in the incidental detection of early-stage renal tumors [3]. Early diagnosis is critical, as renal cancer often remains asymptomatic in its initial stages but can progress rapidly if left untreated. Accurate and timely interpretation of renal CT scans is essential for clinical decision-making, including diagnosis, staging, and treatment planning. However, increasing imaging volumes and inter-observer variability pose challenges for radiologists. As such, the development of intelligent systems to support or automate report generation from renal CT images holds significant potential to enhance diagnostic efficiency, reduce variability, and ultimately improve patient care.
@@
Recent years have witnessed remarkable progress in large model technologies, particularly in the development of large language models (LLMs) and vision-language models (VLMs). These foundation models have brought transformative advances across many domains, including biomedical informatics and computer-aided diagnosis. Models such as GPT-4, PaLM, and Med-PaLM have demonstrated the ability to understand clinical narratives, answer medical questions, and even provide diagnostic suggestions, acting as virtual clinical assistants in various scenarios [4, 5, 6]. Similarly, vision-language models such as BioViL and LLaVA-Med integrate visual and textual modalities to support medical image interpretation, question answering, and image-report alignment [7, 8].
These technological breakthroughs have led to a surge of interest in medical imaging applications, particularly in automating radiology workflows. One rapidly growing area is radiology report generation, where deep models learn to produce textual reports directly from medical images. Most existing work in this space has focused on chest X-rays, leveraging large-scale public datasets such as MIMIC-CXR [9] and CheXpert [10], and achieving promising results through image captioning, retrieval-augmented generation, and transformer-based encoder-decoder frameworks [11].
However, relatively little attention has been paid to radiology report generation for other anatomical regions. In particular, there is a lack of work focusing on renal imaging. Existing studies on kidney CT or MRI mainly emphasize classification tasks, such as tumor detection or subtype prediction [12, 13, 14, 15], rather than comprehensive report generation. This gap is partly due to several challenges unique to renal radiology.
First, there is a scarcity of large-scale, publicly available renal CT datasets with paired expert reports, which limits the development and benchmarking of generative models in this domain. Second, unlike chest imaging, there is no widely accepted or standardized reporting template for renal CT scans, especially in the context of renal cancer, making the target output space more diverse and ambiguous. Third, evaluating the quality of generated radiology reports remains a fundamental challenge. Traditional natural language processing (NLP) metrics such as BLEU, ROUGE, and METEOR focus on surface-level lexical overlap and fail to capture clinically critical information. In medical contexts, a single word or numerical value can determine the diagnostic conclusion, meaning that minor errors may have major clinical implications [16].
In this work, we aim to bridge this gap by exploring the feasibility of automated renal radiology report generation. We collect real-world clinical data from the University of Florida (UF) Health System, and, guided by expert radiologists and nephrologists, construct a structured reporting schema tailored for renal imaging. We propose a two-step generation pipeline: the first stage detects key radiological features from 2D CT slices, and the second stage integrates these findings to generate coherent and clinically meaningful textual reports. To the best of our knowledge, this study represents one of the first systematic efforts toward generative modeling in kidney imaging, and lays foundational work for future kidney-specific AI applications in radiology.
2 Statement of Significance
Problem
Radiology report generation has primarily focused on chest X-rays and high-level classification tasks, with limited exploration of generating structured reports for renal cancer using CT slices.
What is Already Known
Recent advancements in computer vision (CV) and large language models (LLMs) have enabled vision-language models (VLMs) to support a wide range of medical imaging tasks, including radiology report generation. These systems have demonstrated success in generating descriptive reports for chest X-rays. However, to date, no system has been specifically designed to generate reports for renal imaging, particularly in the context of kidney cancer.
What This Paper Adds
@@ 2 Statement of Significance
We present the first system for renal CT report generation using a two-stage pipeline. Based on real-world UF Health data, our method extracts abnormality features and generates clinically meaningful reports.
Who Would Benefit from the New Knowledge in This Paper
AI researchers can leverage this work as a foundation for extending radiology report generation to less-explored anatomical regions. Radiologists and nephrologists may benefit from improved report standardization, enhanced diagnostic support, and reduced manual workload in renal cancer care.
3 Related Work
@@ 3.1 Renal Abnormality Diagnosis
Several studies have demonstrated the effectiveness of deep learning approaches in diagnosis kidney abnormalities from CT scans. Alzu’bi et al. [12] proposed a method for classifying renal tumors, cysts, and stones from 2D CT slices using a new dataset. Similarly, Islam et al. [13] developed vision transformer and transfer learning models to differentiate between normal kidneys and those with tumors, cysts, or stones, also based on 2D CT images. Yildirim et al. [15] focused specifically on the detection of kidney stones using coronal CT slices, employing a deep learning architecture for automatic identification. Uhm et al. [14] addressed a more complex diagnostic task by classifying kidney cancer into five subtypes using multi-phase 3D CT scans.
Although these methods demonstrate promising diagnostic capabilities, they primarily focus on generating high-level classification results. None of these works provide a detailed characterization of abnormality features or generate comprehensive radiology reports that mimic clinical narratives. Our work aims to bridge this gap by incorporating structured abnormality features and producing interpretable, report-style outputs.
@@ 3.2 Large Language Models
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation, leading to growing interest in their application to biomedical domains. OpenAI’s GPT-4 has been shown to perform competitively on medical licensing exams and clinical question-answering tasks [17]. Similarly, Meta’s LLaMA series, particularly LLaMA-2 and its derivatives such as MedLLaMA, have been adapted for biomedical tasks through domain-specific pretraining or instruction tuning [18]. Qwen, a large-scale Chinese open-source foundation model developed by Alibaba, has also released variants with capabilities in medical understanding, contributing to the multilingual extension of biomedical LLMs [19]. Beyond general-purpose LLMs, several models have been tailored for healthcare, such as BioGPT [20], trained on PubMed abstracts, and PMC-LLaMA [21], which combines biomedical knowledge with LLaMA’s architecture. These models have been applied to tasks including medical report summarization, entity extraction, and clinical question answering. However, their use in radiology report generation—particularly for complex imaging modalities such as CT scans—remains limited and underexplored.
@@ 3.3 Radiology Report Generation
Radiology report generation has been extensively studied in the context of chest X-ray imaging. Two widely used datasets in this domain are MIMIC-CXR [9] and IU X-Ray [22], both of which provide paired 2D chest X-ray images (including at least posteroanterior and lateral views) and corresponding free-text reports. However, these datasets lack detailed annotations of abnormality features, which limits their utility for structured or explainable report generation. Hamamci et al. [23] introduced a large-scale 3D chest CT dataset that supports multiple tasks, including relevant case retrieval, multi-abnormality detection, and radiology report generation. In a different domain, Lei et al. [24] released a brain MRI dataset that contains both segmentation masks of abnormal regions and manually written radiology reports, enabling grounded report generation with spatial correspondence between images and text.
To the best of our knowledge, our work is the first to focus on radiology report generation for renal imaging. We aim to provide a detailed characterization of abnormality features along with free-text reports, using 2D CT slices of the kidney.
@@ 4.1 Participants
This study was conducted in compliance with the Declaration of Helsinki, and the requirement for informed written consent was waived by the Institutional Review Board (IRB approval number: IRB202100401). The data used in this study were obtained from the UF Integrated Data Repository, which contains CT scans and clinical reports from multiple hospitals within the UF Health system. [^1] https://idr.ufhealth.org/
Participants were identified through a multi-stage filtering process based on narrative radiology reports and corresponding CT scans. First, all available reports were extracted and filtered using renal-related CPT-9 codes (74160, 74170, 74175, 74177, 74178), which represent abdominal and pelvic CT scans performed with or without intravenous contrast. Reports matching these codes were considered to be associated with renal imaging. Next, these reports were matched to corresponding CT scan records. Reports that could not be linked to an existing CT study or whose scan data were unavailable were excluded. Among the remaining report–scan pairs, we selected those that explicitly referenced a specific CT slice, typically in the coronal view.
Some reports included multiple references to distinct image slices, each describing a separate abnormality; therefore, a single report could yield multiple annotations. These slice-referenced reports were manually annotated, resulting in 155 annotations from 129 reports. After excluding cases with no reported abnormalities, missing CT series, or CT slices in non-coronal views, we obtained a curated annotation set consisting of 130 annotations from 108 reports across 97 patients. Table 1 summarizes the patient demographics and lesion characteristics of the final cohort. Notably, there is a discrepancy in the distribution of diagnosis labels between the patient-level and lesion-level summaries. The diagnosis labels in the patient demographics are derived from all recorded ICD codes across a patient’s medical history, whereas the lesion-level labels are limited to findings explicitly referenced by a CT slice within the radiology report. As a result, certain diagnoses may be present at the patient level but absent from the lesion-level annotations.
Figure 1: Workflow for filtering and annotating renal CT reports and scans. Reports were filtered based on renal-related CPT-9 codes, scan availability, and explicit slice-level references. Reports referencing non-coronal slices or lacking abnormalities were excluded. Some reports contained multiple slice references, resulting in 130 manual annotations from 97 patients.
Table 1: Summary of Patient and Lesion Characteristics
Patient Demographics (n = 97)
@@
Figure 2: Data preprocessing pipeline. The pipeline includes three main components: (1) identifying renal-related sentences from the medical report; (2) extracting corresponding renal features; and (3) locating and extracting the 2D CT slice that matches the referenced image series in the report.
@@ 4.2 Data Preprocessing
Our dataset consists of radiology reports, structured renal feature labels, and corresponding CT image slices. The overall preprocessing pipeline is illustrated in Figure 2.
@@ 4.2.1 Report
The original radiology reports contain sections such as clinical history, exam details, and findings across multiple organs. For this study, we focused exclusively on renal-related findings that describe CT-based observations. Because our analysis is based on 2D CT slices, we extracted only sentences that explicitly reference specific CT slice locations (e.g., “image 45 series 4”). These sentences were identified using the Qwen 2.5 large language model [25] with a tailored prompt, followed by manual review to ensure accuracy and completeness. Sentences without explicit image references were excluded, even if they contained potentially relevant renal content. We also excluded slice-referenced sentences describing findings unrelated to renal cancer progression (e.g., nephrolithiasis or hydronephrosis). This filtering ensured that each retained sentence could be aligned with a specific CT slice and remained within the clinical scope of renal cancer. However, this also limited the diversity of renal pathologies represented in the dataset. The prompt used for sentence extraction is provided in Appendix A.
@@ 4.2.2 Feature Labels
Based on guidance from renal specialists and reference materials such as the Radiology Assistant [26] and RadReport [27], we extracted eight renal feature labels: position, size, exophytic, attenuation, enhancement, cyst, mass, and tumor. We followed the annotation schema from [19], using a large language model to extract and structure these labels, followed by manual verification. In cases where the lesion type (cyst, mass, or tumor) could not be directly determined from the report, nephrology experts assisted in the inference. When certain features were not explicitly mentioned in the report, the corresponding feature labels were marked as Unknown. For qualitative size descriptions such as “subcentimeter,” we standardized the value to 0.5 cm to ensure consistent numerical representation across the dataset. Many feature categories—particularly growth pattern and enhancement—were frequently unreported in the original reports, resulting in a substantial number of Unknown entries (see Table 1). We chose not to impute or assign default values, as omission in clinical documentation does not imply absence. Preserving this missingness reflects the inherent ambiguity of real-world reports and ensures that downstream models are trained and evaluated under realistic, incomplete conditions. The full prompt used for feature label extraction is provided in Appendix A.

### v2

# A Clinically Informed Two-Stage Framework for Renal CT Report Generation

@@ Abstract
Objective Renal cancer is a common malignancy and an important cause of cancer-related mortality. Computed tomography (CT) plays a central role in early detection, staging, and treatment planning. The growing volume of CT studies increases radiologists’ workload and requires specialized training to reliably interpret images and document findings. However, automatically generating such reports remains challenging as it requires integrating visual interpretation with clinical reasoning. Advances in artificial intelligence (AI), including large language models (LLMs) and vision-language systems, may help reduce radiologists’ reporting burden, enhance diagnostic accuracy, and ultimately improve patient care and clinical efficiency.
Methods We propose a clinically informed, two-stage framework for automatic radiology report generation from single-slice renal CT images. In the first stage, we detect structured clinical features from each 2D image using a multi-task learning model. In the second stage, we generate free-text reports conditioned on both the input image and the detected clinical features, leveraging a vision-language foundation model. To assess clinical accuracy and consistency, we further extract generated clinical features from model-generated reports and compare them with expert-annotated ground truth. We use an expert-labeled dataset for training and evaluation.
Results Experiments on an expert-labeled dataset show that incorporating detected clinical features improves both the quality and clinical fidelity of the generated reports. The model achieved an average AUC of 0.75 across key imaging features and reached a maximum METEOR score of 0.33 in report generation, indicating enhanced clinical faithfulness and reduced template-driven hallucinations.
Conclusion Coupling structured clinical feature detection with conditioned report generation offers a promising approach to integrate structured prediction with free-text drafting for renal CT reporting, enhancing interpretability and clinical fidelity. The results also highlight the importance of clinically meaningful evaluation metrics when developing medical AI systems.
keywords:
Radiology report generation , Clinical features detection , Computed tomography (CT) , Vision–language models , Explainable AI in radiology
journal: Journal of Biomedical Informatics
@@ 1 Introduction
Renal cancer, also known as kidney cancer, is a malignancy with increasing global incidence. Each year, an estimated 400,000 new cases are diagnosed, resulting in approximately 175,000 deaths worldwide [1]. Projections suggest that incidence rates will continue to rise over the next decade, underscoring the public health significance of this disease. In the United States alone, more than 81,000 new cases and over 14,000 deaths are expected in 2024 [2]. Although mortality rates have declined by 1–2% annually due to advances in treatment, overall incidence continues to increase [2]. The widespread adoption of cross-sectional imaging—particularly computed tomography (CT)—has led to increased incidental detection of early-stage renal tumors [3]. Early diagnosis is critical, as renal cancer is often asymptomatic in its initial stages but can progress rapidly if untreated. Accurate and timely interpretation of renal CT scans is essential for clinical decision-making, including diagnosis, staging, and treatment planning. However, the growing volume of imaging studies, coupled with inter-observer variability among radiologists, poses significant challenges. These limitations highlight the need for intelligent systems capable of supporting or augmenting radiology report generation from renal CT images, which may improve diagnostic efficiency, reducing variability, and ultimately enhance patient care.
@@
Recent advances in large model technologies, including large language models (LLMs) and vision-language models (VLMs), have transformed biomedical informatics and computer-aided diagnosis. Models such as GPT-4 [4], PaLM [5], and Med-PaLM [6] have demonstrated the ability to understand clinical narratives, answer medical questions, and even provide diagnostic suggestions, functioning as virtual clinical assistants in various scenarios [6, 7, 8]. Similarly, vision-language models (VLMs) such as BioViL [9] and LLaVA-Med [10] integrate visual and textual modalities to support medical image interpretation, question answering, and image-report alignment [11, 12]. These technological breakthroughs have led to a surge of interest in medical imaging applications, particularly in automating radiology workflows. One rapidly growing area is radiology report generation, where deep models learn to produce textual reports directly from medical images. Most existing work in this space has focused on chest X-rays, leveraging large-scale public datasets such as MIMIC-CXR [13] and CheXpert [14].
Despite growing interest in radiology report generation, relatively little attention has been given to anatomical regions beyond the chest. In particular, renal imaging remains an underexplored area. Most existing studies on kidney CT or MRI primarily focus on classification tasks, such as tumor detection or subtype prediction, rather than comprehensive report generation [15, 16, 17, 18]. Several challenges contribute to this gap. First, there is a lack of large-scale, publicly available renal CT datasets with paired expert-authored reports, limiting model development and evaluation. Second, unlike chest imaging, renal CT lacks a widely accepted or standardized reporting template, resulting in a more diverse and ambiguous target output space. Third, evaluating the quality of generated radiology reports remains inherently difficult. Conventional natural language generation (NLG) metrics—such as BLEU, ROUGE, and METEOR—primarily measure lexical overlap and often fail to capture clinically meaningful content. In clinical practice, even minor discrepancies can substantially alter the diagnostic interpretation [19].
Figure 1: Dataset curation and model framework. (a) From expert renal CT radiology reports we derive structured clinical (abnormality) features—position, size, exophytic growth, attenuation, enhancement, cyst, mass, tumor—and a reference slice index; these align each 2D CT slice with labels. (b) From a 3D CT volume, we retrieve the labeled slice and apply intensity clipping to form the 2D input image. (c) Two-stage pipeline: Stage 1 detects clinical features from the image; Stage 2 conditions an LLM-based report generator on the image and the detected features via a structured prompt to produce a narrative clinical report.
In this study, we explore automated renal radiology report generation using real-world clinical data from the UF Health Integrated Data Repository (IDR). We construct a custom structured schema—guided by expert radiologists—for describing renal CT findings. We introduce a two-stage pipeline: first, key radiological features—referred to as clinical features—are detected from individual 2D CT slices; second, a vision-language model generates free-text reports conditioned on both the detected features and the corresponding CT slice. This modular design enhances interpretability and clinical relevance while enabling separate optimization and evaluation of feature extraction and narrative generation.
To our knowledge, this study represents one of the first systematic efforts in generative modeling for kidney imaging. It establishes a clinically informed framework integrating multimodal learning and domain-specific supervision, laying the groundwork for future kidney-specific AI applications in radiology and related domains.
Statement of Significance
Problem Writing renal CT reports is time-consuming and may omit important details. Automatic solutions are needed to improve efficiency and completeness and ultimately support clinicians.
Already Known Most radiology report generation studies have focused on chest X-rays with promising results. However, applications to renal CT remain underexplored.
@@
This Paper Adds We introduced the first system for renal CT report generation using a clinically informed two-stage pipeline. The method detects structured features and generates coherent reports aligned with expert findings, evaluated on real-world data from UF Health IDR.
Benefit Radiologists and nephrologists can benefit from improved report consistency and reduced documentation burden. AI researchers may use this framework to extend multimodal generation to underexplored anatomical regions.
2 Related Works
@@ 2.1 Renal Abnormality Diagnosis
Deep learning approaches have shown promise in diagnosing kidney abnormalities from CT scans. For example, Alzu’bi et al. [15] proposed a method for classifying renal tumors, cysts, and stones from 2D CT slices using a newly introduced dataset. Similarly, Islam et al. [16] developed vision transformer and transfer learning models to distinguish normal kidneys from those with tumors, cysts, or stones, also based on 2D CT images. Yildirim et al. [18] focused specifically on detecting kidney stones using coronal CT slices, employing a deep learning model for automated identification. Uhm et al. [17] addressed a more complex task by classifying kidney cancer into five subtypes using multi-phase 3D CT scans. While these methods demonstrate promising diagnostic performance, they primarily focus on high-level classification tasks and do not capture detailed radiological findings. None of the above studies aim to generate comprehensive radiology reports that mimic expert-written narratives. In contrast, our work detects fine-grained clinical features and generates interpretable, report-style outputs that closely mimic real-world clinical documentation.
@@ 2.2 Large Language Models
LLMs have shown impressive capabilities in natural language understanding and generation, with growing applications in biomedicine. OpenAI’s GPT-4, for instance, performs competitively on medical licensing exams and clinical question-answering tasks [20]. Meta’s LLaMA series, particularly LLaMA-2 and its biomedical adaptation MedLLaMA, have been fine-tuned for healthcare tasks through domain-specific pretraining and instruction tuning [21]. Qwen, an open-source foundation model developed by Alibaba, has introduced medical variants that contribute to the multilingual expansion of biomedical LLMs [22]. Beyond general-purpose LLMs, several models have been explicitly designed for biomedical use. Examples include BioGPT [23], trained on PubMed abstracts, and PMC-LLaMA [24], which integrates biomedical knowledge into the LLaMA architecture. These models have been applied to tasks such as clinical question answering, entity extraction, and medical report summarization. Despite these advances, the use of LLMs for radiology report generation—especially with complex imaging modalities such as CT scans—remains limited and largely underexplored.
@@ 2.3 Radiology Report Generation
Radiology report generation has been extensively studied in the context of chest X-ray imaging. Two widely used datasets in this domain are MIMIC-CXR [13] and IU X-Ray [25], both of which provide paired 2D chest X-ray images (typically including posteroanterior and lateral views) and corresponding free-text reports. However, these datasets lack detailed annotations of clinical findings, which limits their utility for structured or explainable report generation. Hamamci et al. [26] introduced a large-scale 3D chest CT dataset that supports multiple tasks, including relevant case retrieval, multi-abnormality detection, and radiology report generation. In a different domain, Lei et al. [27] released a brain MRI dataset containing both segmentation masks of abnormal regions and manually written reports, enabling spatially grounded report generation.
@@ 3.1 Participants
The data used in this study were obtained from the UF Health Integrated Data Repository (IDR) [^1] https://idr.ufhealth.org/. This study was approved by the Institutional Review Board (IRB202100401), and the requirement for written informed consent was waived. All procedures were conducted in accordance with the Declaration of Helsinki.
Figure 2 illustrates the full pipeline for filtering renal CT reports and scans. Participants were identified through a multi-stage filtering process based on narrative radiology reports and their corresponding CT scans. Initially, all available renal CT reports between December 2, 2011 and August 24, 2024 were extracted and filtered using renal-related CPT-9 codes (74160, 74170, 74175, 74177, 74178), which correspond to abdominal and pelvic CT scans performed with or without intravenous contrast. Reports matching these codes were considered indicative of renal imaging. These reports were then linked to the corresponding CT scan records. Pairs that could not be matched to an existing CT study or lacked available scan data were excluded. From the remaining report–scan pairs, we selected those explicitly referencing a specific CT slice.
Some reports referenced multiple distinct image slices, each associated with a different abnormality; thus, a single report could yield multiple annotations. These slice-referenced findings were manually annotated, resulting in 155 annotations derived from 129 reports. After excluding entries with no reported abnormalities, missing CT series, or non-coronal slices, we obtained a curated set of 130 annotations from 108 reports across 97 patients.
Table 1 summarizes the patient demographics and lesion characteristics of the final cohort. Note that diagnosis labels differ between the patient level and lesion level. Patient-level diagnoses were extracted from all available ICD codes in each patient’s medical history, while lesion-level annotations are restricted to abnormalities explicitly referenced by individual CT slices. Consequently, some diagnoses may appear at the patient level but be absent from the lesion-level summary.
Figure 2: Workflow for filtering renal CT reports and scans. Reports were filtered based on renal-related CPT-9 codes, scan availability, and explicit references to individual CT slices. Reports referencing non-coronal slices or lacking abnormalities were excluded. Some reports contained multiple slice references, resulting in 130 manual annotations from 97 patients.
Table 1: Summary of patient and lesion characteristics.
Patient characteristics (n = 97)
@@ 3.2 Data Preprocessing
Our dataset comprises radiology reports, structured renal feature labels, and corresponding CT image slices. The data preprocessing steps are summarized in Figure 1(a) and (b).
@@ 3.2.1 Report
The original radiology reports include sections such as clinical history, examination details, and imaging findings across multiple organ systems. For this study, we focused specifically on renal-related findings derived from CT observations. Given that our analysis centers on 2D CT slices, we extracted only sentences explicitly referencing specific image locations (e.g., “series 4 image 167”). To identify these slice-referenced sentences, we employed the Qwen 2.5 large language model [28] with a customized prompt, followed by manual review to ensure precision and completeness, as illustrated in Figure 1(a). Sentences without explicit image references were excluded, even if they contained potentially relevant renal content. Additionally, slice-referenced sentences describing findings unrelated to renal cancer progression (e.g., nephrolithiasis or hydronephrosis) were removed. This filtering ensured that each retained sentence aligned with a specific CT slice and remained within the clinical scope of renal cancer. The full prompt used for sentence extraction is provided in  A.
@@ 3.2.2 Feature Labels
Guided by physicians with sub-speciality expertise in the evaluation of renal tumors and informed by reference materials such as the Radiology Assistant [29] and RadReport [30], we defined eight renal feature labels: position, size, growth pattern (exophytic/endophytic), attenuation, enhancement, cyst, mass, and tumor. Following the annotation protocol described in [28], a large language model was used to extract and structure these labels from relevant sentences, followed by manual verification. When lesion type (e.g., cyst, mass, or tumor) could not be directly inferred, physician experts adjudicated the findings on a case-by-case basis. Feature values not explicitly mentioned were marked as unknown. For qualitative size descriptions (e.g., “subcentimeter”), values were standardized to 0.5 cm for consistency.
Several features—especially growth pattern and enhancement—were frequently omitted in original reports, leading to a high prevalence of Unknown entries (see Table 1). We intentionally chose not to impute or assign default values. Preserving missingness reflects the ambiguity inherent in real-world reports and allows downstream models to be trained and evaluated under realistic conditions. The full prompt used for feature extraction is provided in Appendix A.